6.4.6 Question H Use the method of variation of parameters to determine a particular solution to the given equation. y + 8
Answer #1

9+819 = tard 90) sec (90) integrate both side wat we get y +817 = Sec 90 ,0 <oches Alexilares eq m2 + 8120 m = 797 Cot